What is the freezing point of a solution containing 45.39 g 5-bromo-2-naphthoic acid (C11H7BrO2), a nonelectrolyte, and 160.7 g acetic acid (CH3COOH)? The normal freezing point of acetic acid is 16.60 °C and Kfp for acetic acid is 3.90 °C/m.
Answer #1

Answer

12.21℃

Explanation

Freezing point depression(∆Tf) is calculated by the following formula

∆Tf = Kfp × b × i

where,

Kfp = freezing point depression constant, 3.90℃/m

b = molality of solute

i = Van't Hoff factor of solute , 1

molality is defined as number of moles of solute per liter of solution

number of moles of C11H7BrO2 = 45.39g / 251.08g/mol = 0.1808mol

molality of C11H7BrO2 = ( 0.1808mol/160.7g) ×1000g = 1.125m

∆Tfb = 3.90℃/m × 1.125m × 1

∆Tfp = 4.39℃

Freezing point of the solution = 16.60℃ - 4.39℃ = 12.21℃